General Meeting Summary 7/6/17

From cua_phy
Jump to navigation Jump to search

DETECTORS AT JLAB(Arthur)

  • Clean (TEDF) and semi-clean (EEL 102) rooms
  • Clean room is for Neutral Particle Spectrometer activities, semi-clean room for HMS aerogel Cherenkov detector refurbishing
  • Shelves, tables installed
  • Most hardware for data acquisition system available (VMEX, TI, fADC, NIM modules), still need a few more items, e.g. NIM crate, and guidance on system setup - check with Brad and Steve, also note that there are additional instruments in EEL126
  • Required safety documents for both rooms have been submitted - meeting with B. Manzlak scheduled
  • Submitted request to install lock on door to clean room - follow up with W. Akers
  • Discussion about cutting/polishing equipment
  • SHMS aerogel Cherenkov detector cosmic test setup
  • scintillator was installed on back of detector, but to run need to discuss with GEM collaboration about trigger setup - for GEMs drift chambers and first two planes of scintillators are used
  • SOS aerogel Cherenkov detector
  • located in EEL126 on floor, 5-6 PMTs missing
  • check with D. Mack (original owner J. Napolitano) about removing remaining PMTs for HMS aerogel refurbishment, also if can remove aerogel - refractive index is 1.05 and it is old generation aerogel (not hydrophobic)


PBWO4 STUDIES AND HMS PMTS AT CUA (Richard)

  • Mechanical assembly of crystal prototype done, calibrations ongoing
  • initial results indicate that vibrations are very small
  • setup for load cell calibration has been constructed
  • Optimization of uniformity of crystal light collection
  • Testing impact of crystal surface condition on light uniformity - reflectors, surface properties
  • Also testing BGO crystal characteristics in comparison to PbWO4
  • PbWO4 reference (J24) result from June 2017 is 14 pe/MeV
  • Discussion about double peak in pedestal in light yield measurements
  • possible sources: noise in signal cabling or changing gate, e.g. from discriminator
  • check for noise in signal cable: look on scope, then unplug signal from ADC and run DAQ without signal, check pedestal
  • check for changing gate widths: use regular pulser instead of scintillator for gate and check on scope for possible changes
  • HMS PMTs
  • setup darkbox for testing
  • planning to start removing PMTs from cylinders using solvent
  • possible options for PMT replacements: SOS aerogel Cherenkov, SHMS aerogel Cherenkov/Bates spares


DATA ANALYSES AND SIMULATIONS (Andres)

  • Work on analysis of SHMS aerogel Cherenkov detector data ongoing
  • Simulations for detector configuration for pion/kaon structure function studies ongoing
  • No update on DVCS analysis


DETECTORS AT CERN (Rachel)

  • Material budget comparison between phase 1 and 2 suggests that detector material budget estimation from TDR was too optimistic
  • Actual performance of detector layer 1 does not match design - requires more detailed studies of the impact on physics
  • EPS-HEP meeting ongoing in Venice/Italy including talks on pixel detectors
  • LHC currently resolving a problem that prevents from running


NEXT MEETING: WED. 12 JULY at 10:00 AM (EDT)